OpenStudy (anonymous):

The selling price of a fridge ia $799.70. The markup is 10% of the dealers cost, what is the dealers cost of the fridge?

OpenStudy (anonymous):

x*110% = 799.70 --> x = 799.70 / 1.10 = $727.00

OpenStudy (anonymous):

d( 1+ 10%) = 799.70 ->d = 799.70/ 1.1 = $727